### v2

@@ Abstract
We study decision-making problems where data comprises points from a collection of binary polytopes, capturing aggregate information stemming from various combinatorial selection environments. We propose a nonparametric approach for counterfactual inference in this setting based on a representative agent model, where the available data is viewed as arising from maximizing separable concave utility functions over the respective binary polytopes. Our first contribution is to precisely characterize the selection probabilities representable under this model and show that verifying the consistency of any given aggregated selection dataset reduces to solving a polynomial-sized linear program. Building on this characterization, we develop a nonparametric method for counterfactual prediction. When data is inconsistent with the model, finding a best-fitting approximation for prediction reduces to solving a compact mixed-integer convex program. Numerical experiments based on synthetic data demonstrate the method’s flexibility, predictive accuracy, and strong representational power even under model misspecification.
keywords:
discrete choice, combinatorial optimization, nonparametric modeling, representative agent model
1 Introduction
@@
Many decision-making problems involve agents selecting from complex combinatorial feasible sets, such as paths in networks, assignments in matching markets, or constrained item bundles. Given data on such choices, a central challenge is to infer models that both explain observed behavior and enable prediction in counterfactual environments not previously encountered. For instance, using historical congestion data from an evolving city transportation network, how would the introduction of a new link reshape congestion patterns? Such inferential capabilities are essential for understanding system behavior and for informing downstream network and system design.
[…]
@@
In particular, we consider settings where the observed dataset comprises tuples of the form $\{(\boldsymbol{p}^{(k)},\mathcal{X}^{(k)}):k\in[K]\}$, where $K$ is the total number of historical observations, $\mathcal{X}^{(k)}\subseteq\{0,1\}^{n}$ is a combinatorial selection set, and and the respective $\boldsymbol{p}^{(k)}$ lies in the convex hull $\text{conv}(\mathcal{X}^{(k)}).$ While the set $\mathcal{X}^{(k)}$ may be understood to be capturing the feasible set in a given network (or) a combinatorial selection environment, the respective $\boldsymbol{p}^{(k)}$ can be understood as the resulting selection data aggregated/averaged across individual agents. For instance, the set $\mathcal{X}^{(k)}$ may correspond to the set of directed paths available in a transportation network configuration (captured by the flow conservation constraints) and the respective $\boldsymbol{p}^{(k)}$ may capture the empirical aggregated flow of agents traversing these paths. The feasible set instances, denoted by $\mathcal{X}^{(1)},\ldots,\mathcal{X}^{(K)},$ may differ among themselves due to different collections of variables being forced to be zero at different instances, say, due to link dysfunctions/unavailability in networks (or) unavailable pairings at a given point of time.
@@
How might we model such a dataset $\{(\boldsymbol{p}^{(k)},\mathcal{X}^{(k)}):k\in[K]\},$ capturing aggregated selection statistics in combinatorial settings, so that we can perform counterfactual predictions for selections in a new variant, $\mathcal{X}^{\text{new}},$ differing from all the selection sets $\mathcal{X}^{(1)},\ldots,\mathcal{X}^{(K)}$ observed in the past? The significance of this research question spans diverse settings, as indicated in the table below.
Application
@@
In this work, we propose to respond to such challenges by modeling the selection data $\{(\boldsymbol{p}^{(k)},\mathcal{X}^{(k)}):k\in[K]\}$ via the aforementioned representative agent model $\boldsymbol{p}^{(k)}=\arg\max_{\boldsymbol{x}\in\text{conv}(\mathcal{X}^{(k)})}\{\boldsymbol{\nu}^{\top}\boldsymbol{x}-c(\boldsymbol{x})\},$ for $k\in[K].$ In particular, we focus on the case where $c(\boldsymbol{x})$ is in a separable additive form and we call the resulting model as the Separable Representative Agent Model (abbreviated as S-RAM). Formally, in the S-RAM, the representative agent selects a vector
$\boldsymbol{x}^{*}=\arg\max_{\boldsymbol{x}\in\text{conv}(\mathcal{X})}\left\{\boldsymbol{\nu}^{\top}\boldsymbol{x}-\sum_{j\in[n]}c_{j}(x_{j})\right\},$ (S-RAM)
where $\boldsymbol{\nu}\in\mathbb{R}^{n}$ is a vector of deterministic utilities and $\{c_{j}(x_{j}):j\in[n]\}$ are convex perturbation functions. Besides proposing this model, one of our key contributions is to precisely characterize the collection of $(\boldsymbol{p}^{(k)},\mathcal{X}^{(k)})$ tuples which can be modeled by S-RAM. Such exact characterizations have been critical in discrete choice modeling, both in leading to crisp understanding of the representative abilities of choice models and in performing counterfactual inference over product assortments unseen in the datasets; see, Fudenberg et al. (2015); Ruan et al. (2022). Similarly, our modeling of combinatorial selection data via S-RAM leads towards the following contributions:
@@
Empirical performance. Through experiments with synthetic data on network problems including shortest path, longest path and assignment, we demonstrate the practical flexibility, predictive accuracy, and strong representational power even under model misspecification of our framework.
@@
Figure 1: A new framework for counterfactual prediction with a summary of main contributions
2 Problem Statement and Related Literature
Problem Statement. Consider $\mathcal{X}:=\{\boldsymbol{x}\in\{0,1\}^{n}\mid\boldsymbol{A}\boldsymbol{x}\leq\boldsymbol{b}\},$ with $\boldsymbol{A}\in\mathbb{R}^{m\times n}$ and $\boldsymbol{b}\in\mathbb{R}^{m}$ characterizing the combinatorial problem of interest. Let $K$ be the number of historical observations and $\mathcal{K}=\{1,\ldots,K\}$. Suppose for any $k\in\mathcal{K}$, we observe $\boldsymbol{p}^{(k)}\in\text{conv}(\mathcal{X}^{(k)})$ where $\mathcal{X}^{(k)}\subseteq\mathcal{X}$ and and $\boldsymbol{b}\in\mathbb{R}^{m}$. Specifically,
@@ 2 Problem Statement and Related Literature
where $\emptyset\subset\mathcal{I}^{(k)}\subseteq[n]$. Thus, the set $\{\text{conv}(\mathcal{X}^{(k)}):k\in\mathcal{K}\}$ is the collection of historical 010-1 polytopes. Let $\mathcal{P}^{\mathcal{K}}=\{\boldsymbol{p}^{(k)}:k\in\mathcal{K}\}$ be the collection of historical data and $\mathcal{I}^{\mathcal{K}}=\{\mathcal{I}^{(k)}:k\in\mathcal{K}\}$. Given any new structured polytope $\text{conv}(\mathcal{X}^{\text{new}}),$ where $\mathcal{X}^{\text{new}}:=\{\boldsymbol{x}\in\{0,1\}^{n}\mid\boldsymbol{A}\boldsymbol{x}\leq\boldsymbol{b},x_{j}=0,\forall j\notin\mathcal{I}^{\text{new}}\}$, with no prior data, i.e., $\mathcal{I}^{\text{new}}\notin\mathcal{I}^{\mathcal{K}},$ our goal is to: (1) verify whether $\mathcal{P}^{\mathcal{K}}$ is consistent with the S-RAM model hypothesis, and (2) develop a nonparametric data-driven approach for making predictions for the polytope $\text{conv}(\mathcal{X}^{\text{new}})$ that is unseen from the historical data.
Assumption 1
We assume $c_{j}:[0,1]\to\mathbb{R}$ for $j\in[n]$ are univariate, strictly convex and continuously differentiable perturbation functions.
@@
Assumption 2
We assume that
$\text{conv}(\mathcal{X})=\{\boldsymbol{x}\in\mathbb{R}^{n}\mid\boldsymbol{A}\boldsymbol{x}+\boldsymbol{s}=\boldsymbol{b},\boldsymbol{0}\leq\boldsymbol{x}\leq\boldsymbol{e},\boldsymbol{s}\geq\boldsymbol{0}\}.$
@@
Thus, the set $\{\text{conv}(\mathcal{X}^{(k)}):k\in\mathcal{K}\}$ is the collection of historical 010-1 polytopes. Note that all the examples in the table in Section 1 satisfy Assumption 2. For $k\in\mathcal{K},$ the S-RAM solves the following convex program to maximize the expected utility:
$\max_{\boldsymbol{x}}\left\{\boldsymbol{\nu}^{\top}\boldsymbol{x}-\sum_{j\in\mathcal{I}^{(k)}}c_{j}(x_{j})\middle|\ \sum_{j\in\mathcal{I}^{(k)}}a_{ij}x_{j}+s_{i}=b_{i}\forall i\in[m],\ 0\leq x_{j}\leq 1\forall j\in\mathcal{I}^{(k)},s_{i}\geq 0\forall i\in[m]\right\}.$ (1)
@@
$\begin{aligned} \nu_{j}-c_{j}^{\prime}(x_{j}^{*})+\sum_{i\in[m]}\lambda_{i}^{k}a_{ij}+\mu_{j}^{k}-w_{j}^{k}&=0,\quad\forall j\in\mathcal{I}^{(k)},\\ \mu_{j}^{k}x_{j}^{*}=0,\ w_{j}^{k}(1-x_{j}^{*})&=0,\quad\forall j\in\mathcal{I}^{(k)},\\ \sum_{j\in\mathcal{I}^{(k)}}a_{ij}x_{j}^{*}+s_{i}^{*}-b_{i}=0,\ \alpha_{i}s_{i}^{*}&=0,\quad\forall i\in[m],\\ 0\leq x_{j}^{*}\leq 1,\ \mu_{j}^{k},w_{j}^{k}&\geq 0,\quad\forall j\in\mathcal{I}^{(k)},\\ s_{i}^{*}\geq 0,\alpha_{i}&\geq 0,\quad\forall i\in[m],\end{aligned}$ (2)
@@
For each observation $k$, since we observe that $\boldsymbol{p}^{(k)}\in\text{conv}(\mathcal{X}^{(k)})$, we can compute $\bar{s}_{i}=b_{i}-\sum_{j\in\mathcal{I}^{(k)}}a_{ij}p_{j}^{(k)}$, for $i\in[m].$ To obtain the first goal, we aim to verify if $(\boldsymbol{p}^{(k)},\bar{\boldsymbol{s}}^{(k)})$ satisfies the conditions in (2) for all $k\in\mathcal{K}$.
@@
Related Literature. Our paper is closely connected to the machine learning literature on learning and inference in combinatorial optimization, and as well as to the choice modeling literature on combinatorial selection problems.
@@
Learning and Inference in Combinatorial Optimization. Effective structured learning and inference over discrete combinatorial models is challenging. A prominent line of work uses perturbation methods, including perturbed optimizers (Berthet et al., 2020), perturbed structural prediction (Indelman and Hazan, 2021), and stochastic softmax tricks for combinatorial spaces (Paulus et al., 2020). These methods can be viewed as instances of RAMs with perturbation functions and are typically designed for general binary combinatorial optimization. Our work differs in two aspects: (i) we focus on RAMs with separable perturbation functions in combinatorial problems, and (ii) we study 0–1 polytopes, the convex hulls of combinatorial problems, and develop a nonparametric, end-to-end S-RAM framework that remains amenable via standard integer programming solvers. RAM with separable perturbation functions have been shown to yield good tractability in discrete choice while maintaining expressiveness (Ruan et al., 2022). Another stream combines neural networks with combinatorial optimization, using RAMs to make prediction differentiable (Djolonga and Krause, 2017; Dalle et al., 2022; Wilder et al., 2019). While these methods rely on deep learning models to approximate prediction, our framework leverages the polyhedral structure of 010-1 polytopes and the separable structure of S-RAM for tractable, nonparametric counterfactual learning.
@@
Choice Modeling in Combinatorial Problems. A large body of work applies logit models in the random utility model family to network route choice, starting with the MNL for traffic assignment (Dial, 1971) and later route-based extensions that capture overlap and correlated utilities, including path-size, nested, and recursive logit formulations (Chu, 1989; Cascetta et al., 1996; Bekhor and Prashker, 1999; Ben-Akiva and Bierlaire, 1999; Zhou et al., 2012; Mai et al., 2015, 2017; Mai, 2016). More recently, S-RAM–based models have been introduced to traffic network problems, with the Perturbed Utility Route Choice model and its extensions addressing equilibrium assignment (Fosgerau et al., 2022; Yao et al., 2024; Fosgerau et al., 2024). It has been applied beyond transportation, such as product and ancillary pricing in the airline industry (Liu et al., 2022). While these studies are mainly on route-based models, our framework extends S-RAM to a link-based choice model within networks, while enabling a data-driven, nonparametric approach to counterfactual prediction in more general combinatorial settings without restrictive parametric assumptions.

### v3

@@ Abstract.
We study decision-making problems where data comprises historical observations from a collection of binary polytopes, capturing aggregate information stemming from various combinatorial selection environments. We propose a nonparametric approach for counterfactual inference in this setting based on a representative agent model, where the available data is viewed as arising from maximizing separable concave utility functions over the respective binary polytopes. Our first contribution is to precisely characterize the selection probabilities representable under this model and show that verifying the consistency of any given aggregated selection dataset reduces to solving a polynomial-sized linear program. Building on this characterization, we develop a nonparametric method for counterfactual prediction. When data is inconsistent with the model, finding a best-fitting approximation for prediction reduces to solving a compact mixed-integer convex program. Numerical experiments based on synthetic data demonstrate the method’s flexibility, predictive accuracy, and strong representational power even under model misspecification.
conference: ; ;
1. Introduction
@@
Many decision-making problems involve agents selecting from complex combinatorial feasible sets, such as paths in networks, assignments in matching markets, or constrained item bundles. Given data on such choices, a central challenge is to infer models that both explain observed behavior and enable prediction in counterfactual environments not previously encountered. For instance, using historical congestion data from an evolving city transportation network, can we infer how the introduction of a new link would reshape congestion patterns? Such inferential capabilities are essential for understanding system behavior broadly and for informing downstream network and system design.
This paper builds on the following premise: By modeling observed selection data as outcomes of an implicit optimization problem, the induced structure on agent decisions can serve as a vehicle for predicting behavior in new environments.
An elementary, well-studied example of such modeling arises in capturing consumer choice over different product assortments. In particular, a flexible and well-established framework for modeling discrete choice behavior, grounded in microeconomic theory, is the representative agent model (RAM) or equivalently, the perturbed utility model (Fudenberg et al., 2015; McFadden and Fosgerau, 2012; Allen and Rehbeck, 2019), in which a representative agent chooses among $n$ alternatives on behalf of the entire population. Specifically, a representative agent chooses a consumption (or) choice probability vector $\boldsymbol{p}$ that solves the utility maximization
$\boldsymbol{p}=\arg\max\{\boldsymbol{\nu}^{\top}\boldsymbol{x}-c(\boldsymbol{x}):\boldsymbol{x}\in[0,1]^{n},\mathbf{1}^{\intercal}\boldsymbol{x}=1\};$
here the utility function is a linear function perturbed by subtracting a convex function that rewards the agent for diversifying away from the extremal 010-1 vertices capturing deterministic choices. While typical applications of RAM focus on this discrete choice setting, we study its utility in more challenging combinatorial environments.
@@
In particular, we consider settings where the observed dataset comprises tuples of the form $\{(\mathcal{X}^{(k)},\boldsymbol{p}^{(k)}):k\in[K]\}$, where $K$ is the total number of historical observations, $\mathcal{X}^{(k)}\subseteq\{0,1\}^{n}$ is a combinatorial selection set, and and the respective $\boldsymbol{p}^{(k)}$ lies in the convex hull of $\mathcal{X}^{(k)}.$ While the set $\mathcal{X}^{(k)}$ may be understood to be capturing the feasible set in a given network (or) a combinatorial selection environment, the respective $\boldsymbol{p}^{(k)}$ can be understood as the resulting selection data aggregated/averaged across individual agents. For instance, the set $\mathcal{X}^{(k)}$ may correspond to the set of directed paths available in a transportation network configuration (captured by the flow conservation constraints) and the respective $\boldsymbol{p}^{(k)}$ may capture the empirical aggregated flow of agents traversing these paths. The feasible set instances, denoted by $\mathcal{X}^{(1)},\ldots,\mathcal{X}^{(K)},$ may differ among themselves due to different collections of variables being forced to be zero at different instances, say, due to link dysfunctions/unavailability in networks (or) unavailable pairings at a given point of time.
@@
How might we model such a dataset $\{(\mathcal{X}^{(k)},\boldsymbol{p}^{(k)}):k\in[K]\},$ capturing aggregated selection statistics in combinatorial settings, so that we can perform counterfactual predictions for selections in a new variant, $\mathcal{X}^{\text{new}},$ differing from all the selection sets $\mathcal{X}^{(1)},\ldots,\mathcal{X}^{(K)}$ observed in the past? The significance of this research question spans diverse settings, as indicated in Table 1 below.
Table 1. Example application environments
Application (Problem)
@@
In this work, we propose to respond to such challenges by modeling the selection data $\{(\mathcal{X}^{(k)},\boldsymbol{p}^{(k)}):k\in[K]\}$ via the aforementioned representative agent model $\boldsymbol{p}^{(k)}=\arg\max_{\boldsymbol{x}\in\text{conv}(\mathcal{X}^{(k)})}\{\boldsymbol{\nu}^{\top}\boldsymbol{x}-c(\boldsymbol{x})\},$ for $k\in[K].$ In particular, we focus on the case where $c(\boldsymbol{x})$ is in a separable additive form and we call the resulting model as the Separable Representative Agent Model (abbreviated as s-RAM). Formally, in the s-RAM, the representative agent selects a vector
(s-RAM) $\boldsymbol{x}^{*}=\arg\max_{\boldsymbol{x}\in\text{conv}(\mathcal{X})}\left\{\boldsymbol{\nu}^{\top}\boldsymbol{x}-\sum_{j\in[n]}c_{j}(x_{j})\right\},$
where $\boldsymbol{\nu}\in\mathbb{R}^{n}$ is a vector of deterministic utilities and $\{c_{j}(x_{j}):j\in[n]\}$ are convex perturbation functions. Besides proposing this model, one of our key contributions is to precisely characterize the collection of $(\mathcal{X}^{(k)},\boldsymbol{p}^{(k)})$ tuples which can be modeled by s-RAM. Such exact characterizations have been critical in discrete choice modeling, both in leading to crisp understanding of the representative abilities of choice models and in performing counterfactual inference over product assortments unseen in the datasets; see, (Block and Marschak, 1960; Farias et al., 2013; Jagabathula and Rusmevichientong, 2019; Fudenberg et al., 2015; Ruan et al., 2022). Similarly, our modeling of combinatorial selection data via s-RAM leads towards the following contributions:
  1. (1)
@@
  4. (4)
    Empirical performance. Through experiments on synthetic data from network problems including shortest path, longest path, assignment, and constrained subset selection, we demonstrate the practical flexibility, predictive accuracy, and strong representational power of our framework, even under model misspecification. s
@@
Figure 1. A new framework for counterfactual prediction with a summary of main contributions
The remainder of the paper is organized as follows. Section 2 introduces the problem, the model, and the related literature. Section 3 provides a complete characterization of the data representable by the s-RAM model through polynomial-time verifiable necessary and sufficient conditions. In Sections 4 and 5, we illustrate how these characterizations can be used to generate predictions in counterfactual environments and to estimate the best-fitting model, respectively. Section 6 discusses results from numerical experiments on representability and counterfactual predictions.
2. Model Description and Related Literature
2.1. Problem Description
Consider $\mathcal{X}:=\{\boldsymbol{x}\in\{0,1\}^{n}\mid\boldsymbol{A}\boldsymbol{x}\leq\boldsymbol{b}\},$ with $\boldsymbol{A}\in\mathbb{R}^{m\times n}$ and $\boldsymbol{b}\in\mathbb{R}^{m}$ characterizing the combinatorial problem of interest. Let $K$ be the number of historical observations and $\mathcal{K}=\{1,\ldots,K\}$. Suppose for any $k\in\mathcal{K}$, we observe $\boldsymbol{p}^{(k)}\in\text{conv}(\mathcal{X}^{(k)})$ where $\mathcal{X}^{(k)}\subseteq\mathcal{X}$ and and $\boldsymbol{b}\in\mathbb{R}^{m}$. Specifically,
@@ 2.1. Problem Description
where $\emptyset\subset\mathcal{I}^{(k)}\subseteq[n]$. Thus, the set $\{\text{conv}(\mathcal{X}^{(k)}):k\in\mathcal{K}\}$ is the collection of historical 010-1 polytopes. Let $\mathcal{P}^{\mathcal{K}}=\{\boldsymbol{p}^{(k)}:k\in\mathcal{K}\}$ and $\mathcal{I}^{\mathcal{K}}=\{\mathcal{I}^{(k)}:k\in\mathcal{K}\}$ capture the collection of historical observations.
Given any new structured polytope $\text{conv}(\mathcal{X}^{\text{new}}),$ where $\mathcal{X}^{\text{new}}:=\{\boldsymbol{x}\in\{0,1\}^{n}\mid\boldsymbol{A}\boldsymbol{x}\leq\boldsymbol{b},x_{j}=0,\forall j\notin\mathcal{I}^{\text{new}}\}$, with no prior data, i.e., $\mathcal{I}^{\text{new}}\notin\mathcal{I}^{\mathcal{K}},$ our goal is to:
  1. (1)
    verify whether $\mathcal{P}^{\mathcal{K}}$ is consistent with the s-RAM model hypothesis, and
  2. (2)
    develop a nonparametric data-driven approach for making predictions for the polytope $\text{conv}(\mathcal{X}^{\text{new}})$ that is unseen from the historical data.
2.2. The s-RAM model
Assumption 1.
The functions $c_{j}:[0,1]\to\mathbb{R}$ for $j\in[n]$ are univariate, strictly convex and continuously differentiable perturbation functions.
@@ 2.2. The s-RAM model
Assumption 2.
The convex hull of $\mathcal{X}$ $\text{conv}(\mathcal{X})=\{\boldsymbol{x}\in\mathbb{R}^{n}\mid\boldsymbol{A}\boldsymbol{x}+\boldsymbol{s}=\boldsymbol{b},\boldsymbol{0}\leq\boldsymbol{x}\leq\boldsymbol{e},\boldsymbol{s}\geq\boldsymbol{0}\}.$
@@
Thus, the set $\{\text{conv}(\mathcal{X}^{(k)}):k\in\mathcal{K}\}$ is the collection of historical 010-1 polytopes. Note that all the examples in the table in Section 1 satisfy Assumption 2. For $k\in\mathcal{K},$ the s-RAM solves the convex program $\max_{\boldsymbol{x}\in\text{conv}(\mathcal{X}^{(k)})}\big\{\boldsymbol{\nu}^{\top}\boldsymbol{x}-\sum_{j\in\mathcal{I}^{(k)}}c_{j}(x_{j})\big\}$ to maximize the additively perturbed expected utility, which in turn is equivalent to
(1) $\max_{\boldsymbol{x}}\left\{\boldsymbol{\nu}^{\top}\boldsymbol{x}-\sum_{j\in\mathcal{I}^{(k)}}c_{j}(x_{j})\middle|\ \ \begin{aligned} &\sum_{j\in\mathcal{I}^{(k)}}a_{ij}x_{j}+s_{i}=b_{i}\quad\forall i\in[m],\\ &\quad 0\leq x_{j}\leq 1\quad\forall j\in\mathcal{I}^{(k)},\\ &\quad s_{i}\geq 0\quad\forall i\in[m]\end{aligned}\right\}.$
@@
(2) $\begin{aligned} \nu_{j}-c_{j}^{\prime}(x_{j}^{*})+\sum_{i\in[m]}\lambda_{i}^{k}a_{ij}+\mu_{j}^{k}-w_{j}^{k}&=0,\quad\forall j\in\mathcal{I}^{(k)},\\ \mu_{j}^{k}x_{j}^{*}=0,\ w_{j}^{k}(1-x_{j}^{*})&=0,\quad\forall j\in\mathcal{I}^{(k)},\\ \sum_{j\in\mathcal{I}^{(k)}}a_{ij}x_{j}^{*}+s_{i}^{*}-b_{i}=0,\ \alpha_{i}s_{i}^{*}&=0,\quad\forall i\in[m],\\ 0\leq x_{j}^{*}\leq 1,\ \mu_{j}^{k},w_{j}^{k}&\geq 0,\quad\forall j\in\mathcal{I}^{(k)},\\ s_{i}^{*}\geq 0,\alpha_{i}&\geq 0,\quad\forall i\in[m],\end{aligned}$
@@
For each observation $k$, since we observe that $\boldsymbol{p}^{(k)}\in\text{conv}(\mathcal{X}^{(k)})$, we can compute $\bar{s}_{i}=b_{i}-\sum_{j\in\mathcal{I}^{(k)}}a_{ij}p_{j}^{(k)}$, for $i\in[m].$ Then the first goal described in Section 2.1 above translates to verifying if $(\boldsymbol{p}^{(k)},\bar{\boldsymbol{s}}^{(k)})$ satisfies the conditions in (2) for all $k\in\mathcal{K}$.
2.3. Related Literature
@@
Our paper is closely connected to the machine learning literature on learning and inference in combinatorial optimization, and as well as to the choice modeling literature on combinatorial selection problems.
2.3.1. Learning and Inference in Combinatorial Optimization.
@@
Effective structured learning and inference over discrete combinatorial models is challenging. A prominent line of work uses perturbation methods, including perturbed optimizers (Berthet et al., 2020), perturbed structural prediction (Indelman and Hazan, 2021), and stochastic softmax tricks for combinatorial spaces (Paulus et al., 2020). These methods can be viewed as instances of RAMs with perturbation functions and are typically designed for general binary combinatorial optimization. Our work differs in two aspects: (i) we focus on RAMs with separable perturbation functions in combinatorial problems, and (ii) we study 0–1 polytopes, the convex hulls of combinatorial problems, and develop a nonparametric, end-to-end s-RAM framework that remains amenable via standard integer programming solvers. RAM with separable perturbation functions have been shown to yield good tractability in discrete choice while maintaining expressiveness (Ruan et al., 2022).
Another stream combines neural networks with combinatorial optimization, using RAMs to make prediction differentiable (Djolonga and Krause, 2017; Dalle et al., 2022; Wilder et al., 2019). While these methods rely on deep learning models to approximate prediction, our framework leverages the polyhedral structure of 010-1 polytopes and the separable structure of s-RAM for tractable, nonparametric counterfactual learning.
2.3.2. Choice Modeling in Combinatorial Problems.
@@
A large body of work applies logit models in the random utility model family to network route choice, starting with the Multinomial Logit (MNL) model for traffic assignment (Dial, 1971) and later route-based extensions that capture overlap and correlated utilities, including path-size, nested, and recursive logit formulations (Chu, 1989; Cascetta et al., 1996; Bekhor and Prashker, 1999; Ben-Akiva and Bierlaire, 1999; Zhou et al., 2012; Mai et al., 2015, 2017; Mai, 2016). More recently, s-RAM–based models have been introduced to traffic network problems, with the Perturbed Utility Route Choice model and its extensions addressing equilibrium assignment (Fosgerau et al., 2022; Yao et al., 2024; Fosgerau et al., 2024). It has been applied beyond transportation, such as product and ancillary pricing in the airline industry (Liu et al., 2022).
While these studies are mainly on route-based models, our framework extends s-RAM to a link-based choice model within networks, while enabling a data-driven, nonparametric approach to counterfactual prediction in more general combinatorial settings without restrictive parametric assumptions.
2.3.3. Random Utility Interpretation of s-RAM